OpenStudy (anonymous):

Write ans Equation parallel to the given equation through the ordered pair y=4x+6;(1,-3).Please Show Steps ,Instructions Thank you...

OpenStudy (anonymous):

y = 4x-3 So that Would be the answer ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

gud explaining by the way ;p

OpenStudy (anonymous):

@algebra21 this is the way to find the equation of a line parellel to an equation y=4x+6 as i said earlier, the gradient is same...now theres sumthng to do with the point \[m = \frac{y-y_{1}}{x-x_{1}}\] i think u know this \[m(x-x_{1})=y-y_{1}\rightarrow\] u should know this also so u can use this equation to find the equation f a line parellel to y=4x+6 \[y-(-3) = m(x-1)\] \[y+3=4(x-1)\] \[y=4x-3-4\] so the equation is y=4x-7 sorry for the previous mistake, as i thought the x coordinate was "0" remember the gradient formula and u can do anythng :)